OpenStudy (anonymous):

Using systems of linear equations help me!! : How can $12,600 be split between two investments, one paying 6% annually, so that the amounts of interest from the two investments are equal?

OpenStudy (tkhunny):

I'm guessing you'll have toe supply the other investment rate or the total investment rate.

OpenStudy (anonymous):

How would i set it up?

OpenStudy (tkhunny):

You are not understanding my objection. You set this one up by first supplying the entire problem statement. There is insufficient information.

OpenStudy (anonymous):

that's all it says lol -_-

OpenStudy (tkhunny):

Well, then put half the fund into each of two investments both earning 6% and we're done. Silly problem - as written.

OpenStudy (anonymous):

dang i meesed up on wrting its 6% annually and 8% annually

OpenStudy (anonymous):

one paying 6% annually nad one paying 8% annually

OpenStudy (tkhunny):

That's what I was hoping for. Name your two investments. Seriously, this is a very important FORMAL step.

OpenStudy (anonymous):

Using systems of linear equations help me!! : How can $12,600 be split between two investments, one paying 6% annually and one paying 8% annually, so that the amounts of interest from the two investments are equal?

OpenStudy (tkhunny):

Okay, we have that. Now, name your two investments.

OpenStudy (anonymous):

what do you mean lol

OpenStudy (anonymous):

x+y=12,600?

OpenStudy (anonymous):

6x+8y=12600?

OpenStudy (tkhunny):

I mean that if you REALLY want to get it, you must get a little more formal about it. It doesn't take much, but it will save you over and over. x = The Investment @ 6% y = The investment @ 8% Write it down clearly and simply. x + y = 12600 -- You got this one The other one, not so much. "the amounts of interest from the two investments are equal" Calculate the amounts form interest and set them equal. x*0.06 = y*0.08 Does that make sense?

OpenStudy (anonymous):

how would i calculate them though?

OpenStudy (tkhunny):

It's a linear system. You must use both equations to create the correct answers (simultaneous solutions). How do you propose we do that? There are about ten ways to do it. Which have you learned to solve linear systems?

OpenStudy (tkhunny):

The interest from x is x*0.06 - There, I calculated it. The interest from y is x*0.08 - There, I calculated it. I may not mean find an explicit number. Any representation will do.

OpenStudy (anonymous):

substitution

OpenStudy (tkhunny):

Perfect. Let's see it.

OpenStudy (anonymous):

x+y=12,600 y=0.08 then just substitute y in right? and figure out x?

OpenStudy (anonymous):

still dn't get it need full answer

OpenStudy (tkhunny):

Where did you get y = 0.08. This is where that first step will save you. 0.08 is an investment RATE!!! Now, go read the convenient definition that we wrote down. "y = The investment @ 8%" Oh, I see! 'y' is an amount of money, not a rate!!! Why would we EVER write "y = .08". That could not ever makes sense. Adding the investment amounts x + y = 12600 Both amounts of money are equal to another amount of money. "the amounts of interest from the two investments are equal" x*0.06 = y*0.08 And this leaves us with these two beautiful equations. x + y = 12600 x*0.06 = y*0.08 There is our linear system. Solve that.